OpenStudy (anonymous):

Tony's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs $4.60 Tony per pound, and type B coffee costs $5.70 per pound. This month, Tony made 140 pounds of the blend, for a total cost of $706.70 . How many pounds of type A coffee did he use?

OpenStudy (anonymous):

@Hero can you help me out please

OpenStudy (anonymous):

@Hero do know how to set the problem up? then I solve for the answer

hero (hero):

You have to setup two equations, one for the total pounds of coffee, the other for the total cost of coffee per pound. Let A = Pounds of type A coffee B = Pounds of type B coffee Total Pounds of Coffee A + B = 140 Total Cost of Coffee Per Pound 4.60A + 5.70B = 706.70

hero (hero):

Ultimately, you have a system of equations that you'll need to solve: A + B = 140 4.60A + 5.70B = 706.70

OpenStudy (anonymous):

okay i'm lost explain it @Hero

hero (hero):

Let's do it this way. I'll re-write the question. See if that helps.

hero (hero):

Tony's Coffee Shop makes a blend that is a mixture of two types of coffee: Type A and Type B. Tony Blended both types together and made 140 pounds of coffee. Type A coffee costs $4.60 Tony per pound, and type B coffee costs $5.70 per pound for a total cost of 706.70. How many pounds of type A coffee did he use?

OpenStudy (anonymous):

@hero I still don get

hero (hero):

What about it are you confused about?

OpenStudy (anonymous):

i'm asking can you put it in a equation for me then I slove for it @Hero

OpenStudy (anonymous):

@kropot72

hero (hero):

Here's the equation to solve. It's a system: A + B = 140 4.60A + 5.70B = 706.70

hero (hero):

What you have to do is use substitution to write it as a single equation. Notice that A = 140 - B

hero (hero):

Which means you can substitute the expression for A into the second equation to get: 4.60(140 - B) + 5.70B = 706.70

hero (hero):

Now solve for B.

OpenStudy (anonymous):

is it 57 @Hero

hero (hero):

Remember, The goal is to find how many pounds of type A.

hero (hero):

B = 57 is correct. Can you figure out the value of A?